shirley invested some money in an account that paid 9% interest and $250 more than that amount at 10% interest. if her total interest was $101, how much was invested at each rate?
Answer by josmiceli(19441)

total interest = (9% * x) + (10% * (250 + x))
.09x + .1(250 + x) = 101
.09x + 25 + .1x = 101
.19x = 76
x = 400
250 + x = 650
check
.09(400) + .1(650) = 101
36 + 65 = 101
101 = 101
checks OK